# A Generalized Back-Door Criterion for Linear Regression

### v6

@@ Abstract
The conditions under which partial regression coefficients admit a causal interpretation have been discussed in various studies. However, a general condition has not yet been fully established for settings in which mediators or nonlinear terms are included among the explanatory variables. To address this issue, this paper unifies Pearl’s single-door and back-door criteria and proposes a new criterion that enables the identification of causal effects not mediated by covariates. It is shown that, for almost all parameter points in the parameter space of linear causal models, this criterion provides a necessary and sufficient condition under which partial regression coefficients have a linear causal interpretation. In addition, this paper provides a quantitative characterization of post-treatment bias while controlling for confounding bias. These results are applicable to linear models with nonlinear terms, and the decomposition of linear and nonlinear causal effects is also discussed.
@@ 1 Introduction
The conditions under which partial regression coefficients admit a causal interpretation have been discussed in various studies. Among them, one of the most refined conditions is arguably the back-door criterion (Pearl, 1998) or conditional ignorability assumption (Rubin, 1974). When either of these conditions and some additional assumptions are satisfied, partial regression coefficients in linear structural equation models (SEMs), developed by Haavelmo (1943), Haavelmo (1943), Goldberger (1972), and Bollen (1989), can be interpreted as total effects.
However, the following three issues remain unresolved: (1) the control of path-specific effects (i.e., specific mediated effects); (2) the elucidation of the mechanism underlying post-treatment bias; and (3) causal interpretation in the presence of nonlinear terms. Issues (1) and (2) are important topics in mediation analysis. Since nonlinear terms can be regarded as mediators (mediating variables), a deeper understanding of how conditioning on mediators affects partial regression coefficients is required in order to capture complex causal relationships.
In relation to issue (1), the single-door criterion (Pearl, 1998), the back-door criterion for multivariate settings (Nandy et al., 2017), and the sequential ignorability assumption (Imai et al., 2010b) provide partial answers; nevertheless, these conditions cannot accomodate certain path-specific effects. For example, consider the following linear regression equation at the population level:
@@
If we are interested in the mediated effect along the path $X\rightarrow M_{2}\rightarrow Y$ in Figure 1, none of the three conditions clearly tells us whether this effect can be identified by the partial regression coefficient $\beta_{YX\mid M_{1}}$. In fact, this coefficient still coincides with the causal effect of interest, implying that the existing conditions can be generalized.
As shown in equation (1.1), mediation analysis often makes use of post-treatment variables to block certain causal paths. When conditioning on a mediator, one might expect to control only for the mediated effect; however, it may also induce unintended bias. For example, a variable such as $M_{2}$ in Figure 1, which is influenced by multiple variables, is referred to as a collider, and it is well known that conditioning on a collider can introduce bias of this kind. Unfortunately, the mechanism underlying post-treatment bias has not yet been fully elucidated.
In regression analysis, nonlinear terms have been included to capture complex relationships. For example, if the data-generating process of $M_{2}$ is given by $M_{2}=XM_{1}$, the directed edges $X\rightarrow M_{2}$ and $M_{1}\rightarrow M_{2}$ in Figure 1 represent nonlinear causal relationships. In such cases, Imai et al. (2010a) and VanderWeele (2013) examine the identification of total causal effects of $X$ on $Y$ using partial regression coefficients. Nevertheless, these result are based on specific structural assumption; thus, it is necessary to discuss the problem in a more generalized setting.
Against this background, we propose a generalized grahical criterion and provide essentially necessary and sufficient condition under which a partial regression coefficient coincides with the causal effect not mediated by other explanatory variables. In addition, we provide a quantitative characterization of post-treatment bias while controlling for confounding bias. This characterization shows that the joint structure of directed paths and back-door paths is the source of this bias, and that there exist structures that cannot be captured by the conventional concept of path blocking. Finally, we confirm that the above results can be applied to nonlinear data-generating processes and discuss a decomposition of linear and nonlinear causal effects.
The remainder of this paper is organized as follows. Section 2 introduces the key terms as preliminaries. Section 3 formulates the problem we addressed and provides the main theorems. This section also discusses the mechanisms of post-treatment bias and application to nonlinear causal models. All proofs are presented in Appendix.
@@ 2.1 Model Setup
Equation (2.1) can be rewritten as follows:
$X_{i}=c_{i}+\sum_{X_{j}\in\bm{PA}(X_{i})}a_{ij}X_{j}+\tilde{u}_{i},\quad X_{i}\in\bm{V},$ (2.2)
where $c_{i}=\,\mathrm{E}[u_{i}]$ and $\tilde{u}_{i}=u_{i}-\,\mathrm{E}[u_{i}]$. Moreover, the following vector forms of (2.1) and (2.2) are useful:
@@
where $\bm{X}=(X_{1},\dots,X_{N})^{\top}$, $\bm{u}=(u_{1},\dots,u_{N})^{\top}$, and $\tilde{\bm{u}}=(\tilde{u}_{1},\dots,\tilde{u}_{N})^{\top}$. Here, $\bm{A}$ is an $N\times N$ matrix, and $\bm{c}$ is an $N\times 1$ vector. We make the following assumptions for (2.3).
@@
Assumption 2.1 ensures that each graph associated with equation (2.1) is an acyclic directed mixed graph (ADMG), discussed by (Richardson, 2003). In particular, assumption (ii) implies that a causal ordering must be specified. For example, in some cases the graph structure can be identified using methods such as LiNGAM (Shimizu et al., 2006). When the graph structure is unknown, it is necessary to conduct the analysis under each plausible structure and compare the results.
@@ 2.2 Graph Terminology and Notations
An ADMG ia a triple $(\bm{V},\bm{E},\bm{B})$, where $\bm{V}$ is a finite set of vertices, $\bm{E}$ is a set of directed edges, and $\bm{B}$ is a set of bidirected edges. For vertices $a,b\in\bm{V}$, if there is a directed edge from $a$ to $b$ (deneted by $a\to b$), then $a$ is called a parent of $b$, and $b$ is called a child of $a$. Two vertices are said to be adjacent if they are connected by either a directed edge or a bidirected edge. A path is a sequence of distinct vertices such that consecutive vertices are adjacent. A directed path from $a$ to $b$ (denoted by $a\to\to b$) is a path along which all edges are directed and oriented from $a$ toward $b$. If there is a directed path from $a$ to $b$, then $a$ is called an ancestor of $b$, and $b$ is called a descendant of $a$. In what follows, the sets of ancestors and descendants of $a$ are denoted by $\bm{AN}(a)$ and $\bm{DE}(a)$, respectively. A back-door path from $a$ to $b$ (denoted by $a\dashleftarrow b$) is a path between $a$ and $b$ that begins with a bidirected edge or a directed edge pointing into $a$. A v-structure is a path of the form $a\rightarrow c\leftarrow b$, $a\rightarrow c\leftrightarrow b$, $a\leftrightarrow c\leftarrow b$, or $a\leftrightarrow c\leftrightarrow b$, and $c$ is said to be a colider in each v-structure.
@@ 2.3 Causal Path Diagram and the Selective-Door Criterion
Consider equation (2.1). Let $\bm{V}$ denote the set of vertices, and let $\bm{E}$ denote the set of directed edges such that, for each $X_{i}\in\bm{V}$, the set $\bm{PA}(X_{i})$ coincides with the parent set of $X_{i}$. In addition, let $\bm{B}$ denote the set of bidirected edges given by
@@
Then, the triple $(\bm{V},\bm{E},\bm{B})$ is called the causal path diagram associated with equation (2.1).
@@
    for any $X_{k}\in\bm{Z}\cap\bm{DE}(X_{j})$ such that there exists a directed path $X_{j}\rightarrow\rightarrow X_{k}$ that is unblocked by $(\bm{Z}\cup\{X_{j}\})\setminus\{X_{k}\}$, every back-door path $X_{k}\dashleftarrow X_{i}$ is blocked by $\bm{Z}\setminus\{X_{k}\}$, and
@@
Figure 2: An example of causal path diagram
In Figure 2, the set $\{M_{2}\}$ only unblocks the directed path $X\rightarrow M_{1}\rightarrow Y$. However, this set does not satisfy the selective-door criterion relative to $(X,Y)$ due to the existence of the sequence of vertices $X\rightarrow M_{1}\rightarrow M_{2}\leftarrow M_{1}\leftarrow Y$. Since this structure is not a path and $M_{2}$ is not a collider, it can be easily overlooked when relying solely on the conventional concept of path blocking.
@@ 2.4 Controlled Total Effects
In the literature on path analysis and structural equation modeling (SEM), the decomposition of causal effects has been extensively discussed and forms the foundation of mediation analysis. In this framework, the direct effect of $X_{j}$ on $X_{i}$ is given by the path coefficient $a_{ij}$, which corresponds to $X_{j}\to X_{i}$. The indirect effect of $X_{j}$ on $X_{i}$ is defined as the sum of products of path coefficients, which corresponds to each directed path $X_{j}\rightarrow\rightarrow X_{i}$ excluding the directed edge. The total effect of $X_{j}$ on $X_{i}$ is the sum of direct and indirect effects of $X_{j}$ on $X_{i}$. See Alwin and Hauser (1975) and Bollen (1987) for details on these causal effects.
To clarify the interpretation of partial regression coefficients, this paper defines the following partial causal effects.
@@
Let $\bm{G}=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with equation (2.1). For a vertex set $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$, consider the subgraph $\bm{G}^{\prime}$ obtained by removing all edges whose terminal vertices belong to $\bm{Z}$. Then, the $\bm{Z}$-controlled total effect of $X_{j}$ on $X_{i}$, denoted by $\tau_{ij\mid do(\bm{Z})}$, is defined as the total effect of $X_{j}$ on $X_{i}$ restricted to the subgraph $\bm{G}^{\prime}$.
The validity of this measure can be ensured by the following alternative expression of the data-generating process, which is derived solely through successive substitution operations in the linear structural equations.
@@
Assume that (ii) in Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with equation (2.1). Then, for any subset $\bm{S}\subset\bm{V}\setminus\{X_{i}\}$, the following holds:
@@
$\tau_{ij\mid do(\bm{S}\setminus\{X_{j}\})}=\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}\mid do(X_{j}=x_{j}),do(\bm{Z}=\bm{z})],$ (2.5)
where $\bm{Z}=\bm{S}\setminus\{X_{j}\}$, and “dodo” denotes the do-operator (Pearl, 2009). Note that the notation involving the do-operator represents probabilistic concepts of the post-intervention distribution, i.e., the distribution obtained by replacing the right-hand side of the corresponding structural equation with a constant.
@@ 3.2 Main Results
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with equation (2.1). For any set $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$ satisfying the selective-door criterion relative to $(X_{j},X_{i})$, the coefficient $\beta_{ij\mid\bm{Z}}$ in equation (3.1) coincides with the $\bm{Z}$-controlled total effect $\tau_{ij\mid do(\bm{Z})}$.
In the case of linear regression, condition (i) of the selective-door criterion requires that if an explanatory variable $X_{k}\in\bm{Z}$ is a post-treatment variable of $X_{j}$, then one of the following conditions must hold: (1) all directed paths from $X_{j}$ to $X_{k}$ are blocked by other explanatory variables; or (2) all back-door paths from $X_{k}$ to $X_{i}$ are blocked by explanatory variables that also include $X_{j}$. As will be shown later, the converse of Theorem 3.1 also holds almost everywhere in the parameter space. That is, the selective-door criterion constitutes an essentially necessary and sufficient condition for giving a linear causal interpretation to the partial regression coefficients.
@@
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with equation (2.1). For any set $\bm{Z}\subset\bm{V}\setminus(\{X_{i},X_{j}\}$ that blocks every path between $X_{j}$ and $X_{i}$. Then, the coefficient $\beta_{ij\mid\bm{Z}}$ in equation (3.1) is equal to 0.
@@
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with equation (2.1). For any set $\bm{Z}\subset\bm{V}\setminus(\{X_{i},X_{j}\}\cup\bm{DE}(X_{i}))$ that blocks every path between $X_{j}$ and $X_{i}$ except for the directed edge $X_{j}\rightarrow X_{i}$. Then, the coefficient $\beta_{ij\mid\bm{Z}}$ in equation (3.1) coincides with the direct effect of $X_{j}$ on $X_{i}$.
@@
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with equation (2.1). For any set $\bm{Z}\subset\bm{V}\setminus(\{X_{i},X_{j}\}\cup\bm{DE}(X_{j}))$ satisfying the back-door criterion relative to $(X_{j},X_{i})$, the coefficient $\beta_{ij\mid\bm{Z}}$ in equation (3.1) coincides with the total effect of $X_{j}$ on $X_{i}$.
These corollaries were originally established in Spirtes et al. (1998) and Pearl (1998). Note that the single-door criterion (adjustment) excludes descendants of the outcome variable. This reason is explained by the mechanism of post-treatment bias discussed in the following subsection.

### v7

@@ Abstract
The conditions under which partial regression coefficients admit a causal interpretation have been discussed in various studies. However, a general condition has not yet been fully established for settings in which mediators or nonlinear terms are included among the explanatory variables. To address this issue, this paper unifies Pearl’s single-door and back-door criteria and proposes a generalized criterion that enables the identification of causal effects not mediated by other explanatoly variables. In other words, for almost all parameter points in the parameter space of linear causal models, this criterion is a necessary and sufficient condition under which partial regression coefficients have a linear causal interpretation. In addition, this paper provides a quantitative characterization of post-treatment bias while controlling for confounding bias. These results are applicable to linear models with nonlinear terms, and the decomposition of linear and nonlinear causal effects is also discussed.
@@ 1 Introduction
The conditions under which partial regression coefficients admit a causal interpretation have been discussed in various studies. Among existing approaches, the back-door criterion (Pearl, 1998) and the conditional ignorability assumption (Rubin, 1974) are often regarded as among the most refined conditions. When either of these conditions and some additional assumptions are satisfied, partial regression coefficients in linear structural equation models (SEMs), as developed by Wright (1921), Haavelmo (1943), Jöreskog (1970), Goldberger (1972), and Bollen (1989), can be interpreted as total effects.
However, the following three issues remain unresolved: (1) the control of specific mediated effects; (2) understanding the mechanism underlying post-treatment bias; and (3) causal interpretation in the presence of nonlinear terms. Issues (1) and (2) are fundamental topics in mediation analysis. Since nonlinear terms can be regarded as mediators, a deeper understanding of how conditioning on mediators affects partial regression coefficients is necessary for capturing complex causal relationships.
In relation to issue (1), the single-door criterion (Pearl, 1998) and the sequential ignorability assumption (Imai et al., 2010b) provide partial answers; nevertheless, these conditions cannot accommodate certain mediated effects along specific causal paths. For example, consider the following linear regression equation at the population level:
@@
If we are interested in the mediated effect along the path $X\rightarrow M_{2}\rightarrow Y$ in Figure 1, neither condition clearly indicates whether this effect can be identified via the partial regression coefficient $\beta_{YX\mid M_{1}}$. As it turns out, this coefficient still coincides with the causal effect of interest, implying that the existing conditions can be generalized.
As shown in (1.1), mediation analysis often makes use of post-treatment variables to block certain causal paths. When conditioning on a mediator, one might expect to control only for the mediated effect; however, it may also induce unintended bias. For example, a variable such as $M_{2}$ in Figure 1, which is influenced by multiple variables, is referred to as a collider, and it is well known that conditioning on a collider can introduce such bias. Nevertheless, the mechanism underlying post-treatment bias has not yet been fully elucidated.
In regression analysis, nonlinear terms are often included to capture complex relationships. For example, if the data-generating process of $M_{2}$ is given by $M_{2}=XM_{1}$, the directed edges $X\rightarrow M_{2}$ and $M_{1}\rightarrow M_{2}$ in Figure 1 represent nonlinear causal relationships. In such cases, Imai et al. (2010a) and VanderWeele (2013) examine the identification of total causal effects of $X$ on $Y$ using partial regression coefficients. However, these results rely on specific structural assumptions; therefore, it is necessary to discuss this problem in a more generalized setting.
Against this background, we propose a generalized graphical criterion and provide, in a generic sense, a necessary and sufficient condition under which a partial regression coefficient coincides with a causal effect not mediated by other explanatory variables. In addition, we provide a quantitative characterization of post-treatment bias while controlling for confounding bias. This characterization shows that the joint structure of directed paths and back-door paths is the source of this bias, and that there exist structures that cannot be captured by the conventional concept of path blocking. Finally, we show that the above results can be applied to nonlinear data-generating processes and discuss a decomposition of linear and nonlinear causal effects.
The remainder of this paper is organized as follows. Section 2 introduces key terms as preliminaries. Section 3 formulates the problem addressed in this paper and presents the main theorems. This section also discusses the mechanisms underlying post-treatment bias and applications to nonlinear causal models. All proofs are presented in the Appendix.
@@ 2.1 Model Setup
model (2.1) can be rewritten as follows:
$X_{i}=c_{i}+\sum_{X_{j}\in\bm{PA}(X_{i})}a_{ij}X_{j}+\tilde{u}_{i},\quad X_{i}\in\bm{V},$
where $c_{i}=\,\mathrm{E}[u_{i}]$ and $\tilde{u}_{i}=u_{i}-\,\mathrm{E}[u_{i}]$. Moreover, the following vector form is useful:
@@
where $\bm{X}=(X_{1},\dots,X_{N})^{\top}$, $\bm{u}=(u_{1},\dots,u_{N})^{\top}$, and $\tilde{\bm{u}}=(\tilde{u}_{1},\dots,\tilde{u}_{N})^{\top}$. Here, $\bm{A}$ is an $N\times N$ matrix, and $\bm{c}$ is an $N\times 1$ vector. We make the following assumptions on (2.2).
@@
Assumption 2.1 ensures that the graph associated with model (2.1) is an acyclic directed mixed graph (ADMG), discussed by (Richardson, 2003). With respect to the coefficient matrix $\bm{A}$, strictly speaking, it is sufficient that $\bm{A}$ can be transformed into the structure described in (ii) by reordering the indices of the random variables in $\bm{V}$. However, note that unless the graph structure is known, graphical criteria are in most cases not applicable. In some cases, the graph structure can be identified using methods such as LiNGAM (Shimizu et al., 2006). When the graph structure is unknown, it is necessary to conduct the analysis under each plausible structure and compare the results.
@@ 2.2 Graph Terminology and Notations
An ADMG is a triple $(\bm{V},\bm{E},\bm{B})$, where $\bm{V}$ is a finite set of vertices, $\bm{E}$ is a set of directed edges that construct no directed cycles, and $\bm{B}$ is a set of bidirected edges. For vertices $a,b\in\bm{V}$, if they are connected by either a directed edge (denoted by $a\rightarrow b$) or a bidirected edge (denoted by $a\leftrightarrow b$), then $a$ and $b$ are said to be adjacent. If there is a directed edge from $a$ to $b$, then $a$ is called a parent of $b$, and $b$ is called a child of $a$. A path is a sequence of distinct vertices such that consecutive vertices are adjacent. A directed path from $a$ to $b$ (denoted by $a\rightarrow\rightarrow b$) is a path along which all edges are directed and oriented from $a$ toward $b$. A directed cycle is a concatenation of a directed path $a\rightarrow\rightarrow b$ and a directed edge $b\rightarrow a$. If there is a directed path from $a$ to $b$, then $a$ is called an ancestor of $b$, and $b$ is called a descendant of $a$. We denote the sets of ancestors and descendants of $a$ by $\bm{AN}(a)$ and $\bm{DE}(a)$, respectively. A back-door path from $a$ to $b$ (denoted by $a\dashleftarrow b$) is a path between $a$ and $b$ that begins with either a bidirected edge or a directed edge pointing into $a$. A v-structure is a path of the form $a\rightarrow c\leftarrow b$, $a\rightarrow c\leftrightarrow b$, $a\leftrightarrow c\leftarrow b$, or $a\leftrightarrow c\leftrightarrow b$, where $c$ is called a collider.
@@ 2.3 Causal Path Diagram and the Selective-Door Criterion
Consider model (2.1). Let $\bm{V}$ denote the set of vertices, and let $\bm{E}$ denote the set of directed edges such that, for each $X_{i}\in\bm{V}$, the set $\bm{PA}(X_{i})$ coincides with the parent set of $X_{i}$. In addition, let $\bm{B}$ denote the set of bidirected edges given by
@@
Then, the triple $(\bm{V},\bm{E},\bm{B})$ is called the causal path diagram associated with model (2.1).
@@
    for any $X_{k}\in\bm{Z}\cap\bm{DE}(X_{j})$ such that there exists a directed path $X_{j}\rightarrow\rightarrow X_{k}$ that is unblocked by $\bm{Z}\setminus\{X_{k}\}$, every back-door path $X_{k}\dashleftarrow X_{i}$ is blocked by $(\bm{Z}\cup\{X_{j}\})\setminus\{X_{k}\}$, and
@@
Figure 2: An example of unintended post-treatment bias induced by a non-collider
In Figure 2, the set $\{M_{2}\}$ only unblocks the directed path $X\rightarrow M_{1}\rightarrow Y$. However, this set does not satisfy the selective-door criterion relative to $(X,Y)$ due to the existence of the sequence of vertices $X\rightarrow M_{1}\rightarrow M_{2}\leftarrow M_{1}\rightarrow Y$. Since this structure is not a path and $M_{2}$ is not a collider, it can be easily overlooked when relying solely on the conventional concept of path blocking.
@@ 2.4 Controlled Total Effects
In the literature on structural equation modeling, the decomposition of causal effects has been extensively discussed and forms the foundation of mediation analysis. In this framework, the direct effect of $X_{j}$ on $X_{i}$ is given by the path coefficient $a_{ij}$, which corresponds to the directed edge $X_{j}\to X_{i}$. The indirect effect of $X_{j}$ on $X_{i}$ is defined as the sum of products of path coefficients along each directed path $X_{j}\rightarrow\rightarrow X_{i}$, excluding the direct edge $X_{j}\to X_{i}$. The total effect of $X_{j}$ on $X_{i}$ is the sum of the direct and indirect effects. See Alwin and Hauser (1975) and Bollen (1987) for details on these causal effects.
To clarify the interpretation of partial regression coefficients, we define the following partial causal effects.
@@
Let $\bm{G}=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). For a vertex set $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$, consider the subgraph $\bm{G}^{\prime}$ obtained by removing all edges whose terminal vertices belong to $\bm{Z}$. Then, the $\bm{Z}$-controlled total effect of $X_{j}$ on $X_{i}$, denoted by $\tau_{ij\mid do(\bm{Z})}$, is defined as the total effect of $X_{j}$ on $X_{i}$ restricted to the subgraph $\bm{G}^{\prime}$.
The $\bm{Z}$-controlled total effect is the sum of all effects not mediated by variables in $\bm{Z}$. The validity of this measure is justified by the following alternative expression of the data-generating process, which is derived solely through successive substitution operations in the linear structural equations.
@@
Assume that (ii) in Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). Then, for any subset $\bm{S}\subset\bm{V}\setminus\{X_{i}\}$, the following holds:
@@
$\tau_{ij\mid do(\bm{Z})}=\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}\mid do(X_{j}=x_{j}),do(\bm{Z}=\bm{z})],$ (2.4)
where $\bm{Z}=\bm{S}\setminus\{X_{j}\}$, and “dodo” denotes the do-operator (Pearl, 2009). Note that the notation involving the do-operator denotes probabilistic concepts of the post-intervention distribution, i.e., the distribution obtained by replacing the right-hand side of the corresponding structural equation with a constant.
@@ 3.2 Main Results
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). For any set $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$ satisfying the selective-door criterion relative to $(X_{j},X_{i})$, the coefficient $\beta_{ij\mid\bm{Z}}$ in (3.1) coincides with the $\bm{Z}$-controlled total effect $\tau_{ij\mid do(\bm{Z})}$.
In the case of linear regression, condition (i) of the selective-door criterion requires that if an explanatory variable $X_{k}\in\bm{Z}$ is a post-treatment variable of $X_{j}$, then one of the following conditions must hold: (1) all directed paths from $X_{j}$ to $X_{k}$ are blocked by other explanatory variables; or (2) all back-door paths from $X_{k}$ to $X_{i}$ are blocked by explanatory variables other than $X_{k}$. As will be shown later, the converse of Theorem 3.1 also holds almost everywhere in the parameter space. That is, in a generic sense, the selective-door criterion constitutes a necessary and sufficient condition for providing a linear causal interpretation of partial regression coefficients.
@@
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). For any set $\bm{Z}\subset\bm{V}\setminus(\{X_{i},X_{j}\}$ that blocks every path between $X_{j}$ and $X_{i}$. Then, the coefficient $\beta_{ij\mid\bm{Z}}$ in (3.1) is equal to 0.
@@
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). For any set $\bm{Z}\subset\bm{V}\setminus(\{X_{i},X_{j}\}\cup\bm{DE}(X_{i}))$ that blocks every path between $X_{j}$ and $X_{i}$ except for the directed edge $X_{j}\rightarrow X_{i}$. Then, the coefficient $\beta_{ij\mid\bm{Z}}$ in (3.1) coincides with the direct effect of $X_{j}$ on $X_{i}$.
@@
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E},\bm{B})$ be the causal path diagram associated with model (2.1). For any set $\bm{Z}\subset\bm{V}\setminus(\{X_{i},X_{j}\}\cup\bm{DE}(X_{j}))$ satisfying the back-door criterion relative to $(X_{j},X_{i})$, the coefficient $\beta_{ij\mid\bm{Z}}$ in (3.1) coincides with the total effect of $X_{j}$ on $X_{i}$.
These corollaries were originally established in Spirtes et al. (1998) and Pearl (1998). Note that the single-door criterion (adjustment) excludes descendants of the outcome variable. The reason for this exclusion is explained by the mechanism of post-treatment bias discussed in the following subsection.